OpenStudy (anonymous):

I'm so confused! I'm in 6th grade, I have a math test tommorow, and I was sick today, I need help with this!

OpenStudy (anonymous):

Go ahead and post

OpenStudy (anonymous):

Find the GCF of8 and 25

OpenStudy (anonymous):

1

OpenStudy (anonymous):

There is no common factor between 8 and 25 besides 1. The only, let alone greatest, is 1.

jimthompson5910 (jim_thompson5910):

Factor 8 into 4*2 then factor 4 into 2*2 So 8 factors to 2*2*2, ie 8 = 2*2*2 Now factor 25 into 5*5 So 25 = 5*5 Notice how 8 and 25 share no factors whatsoever. So this means that the GCF is 1